25) The price of a lake front lot in a certain housing development is $125,000. This represents 113% 25) increase over the price when the development was first created a decade ago. Find the price of the lot when the development was first created. A) $143,678 B) $58.685 C) $113,700 D) $110,619

Please help me solve this thank you!!!
Answer #1

Let the price of the lot when development was first created be x.

The increase in price is 113% = 1.13x

So, Acc . to question

Current price = x+ 1.13x

=>125000 = 2.13x

=> x= 125000/2.13 = 58685

So , the price of a lot a decade ago is 58685 Dollars

So, option B) 58685 is correct

Please upvote. Thank you.